OpenStudy (anonymous):

How do you find the cube roots of -64?

OpenStudy (lalaly):

-4

OpenStudy (saifoo.khan):

Ninj'ad

OpenStudy (saifoo.khan):

damn!

OpenStudy (anonymous):

make it an imaginary number first

OpenStudy (angela210793):

no just -4

OpenStudy (lalaly):

-4*-4*-4=-64

OpenStudy (anonymous):

actually nvm its a cube root

OpenStudy (anonymous):

u dont need to saifoo stop confusing me

OpenStudy (angela210793):

yea so u can have negative no.s under it :)

OpenStudy (saifoo.khan):

\[\Huge -4\]

OpenStudy (saifoo.khan):

lol. My bad. :(

OpenStudy (anonymous):

hi angela

OpenStudy (anonymous):

i would write that as 4i

OpenStudy (angela210793):

@Hahd Hi!

OpenStudy (anonymous):

zbay -64i thats what you get with 4i

OpenStudy (anonymous):

keep it -4

OpenStudy (saifoo.khan):

\[\Huge \color\red {-4}\]

OpenStudy (anonymous):

\[\color{salmon}{\huge{\checkmark}}\]

OpenStudy (anonymous):

actually there should be 3 roots -4 is only one of them???

OpenStudy (angela210793):

3 roots? O.o

OpenStudy (anonymous):

cube roots of -64???

OpenStudy (anonymous):

there are no roots its not an equation someone's sleepy

OpenStudy (saifoo.khan):

\[\Huge \color\green \blacksquare\]

OpenStudy (nilankshi):

this is rite

OpenStudy (saifoo.khan):

\[\Huge \color\yellow \blacksquare\]

OpenStudy (saifoo.khan):

this one matches! :)

OpenStudy (angela210793):

Saif....teach me tht...

OpenStudy (anonymous):

x^3 = -64 x^3 + 64 = 0 (x + 4)(x^2 + 16 -4x) = 0 x = -4 x^2 + 16 -4x=0 \[x = \frac{4 \pm \sqrt{16 - 64}}{2}\]

OpenStudy (anonymous):

now solve them

OpenStudy (anonymous):

you will get your three

OpenStudy (saifoo.khan):

angela, check your facebook.

OpenStudy (anonymous):

thanksss

OpenStudy (anonymous):

ishaan i hate to poinnt it out but ur b value is wrong

OpenStudy (saifoo.khan):

how about this? \[\Huge \bigodot\]

OpenStudy (saifoo.khan):

Punch here^

OpenStudy (anonymous):

are you sure hahd ?

OpenStudy (anonymous):

u put the constant

OpenStudy (anonymous):

oh nvm

OpenStudy (anonymous):

check that again hahd I am sure it's not like that

OpenStudy (anonymous):

lol its 7 in the morning

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

real answer would be -4

OpenStudy (anonymous):

ready to be sensei today imran good morning : )

OpenStudy (anonymous):

but there are three roots so two additional roots are 4(cos 60+ i sin 60) \[2+i\sqrt{3}\] and its conjugate \[2-i\sqrt{3}\]

OpenStudy (anonymous):

morning ishaan